OpenStudy (anonymous):

Two positive numbers have a difference of 8 and a product of 33. What are these numbers?

OpenStudy (johnweldon1993):

Well lets set up the equations \[\large x - y = 8\] \[\large x \times y = 33\] Lets solve that top equation for 'x' by adding y to both sides \[\large x = y + 8\] Substitute that in for 'x' in the second equation \[\large (y + 8)y = 33\] Distribute \[\large y^2 + 8y = 33\] Lets complete the square there \[\large (y + 4)^2 = 49\] Take the square root of both sides \[\large y + 4 = \sqrt{49}\] Simplify \[\large y + 4 = 7\] Subtract 4 from both sides \[\large y = 3\] Now if we know y = 3 we can substitute that into the original equations to solve for 'x' \[\large x - y = 8\] \[\large x \times y = 33\] Lets put that y = 3 in \[\large x - 3 = 8\] Add 3 to both sides \[\large x = 11\] Now lets check with that second equation \[\large 3 \times 11 = 33 \checkmark \]